### v1

@@ 2.4.3 Ablation studies
The second ablation study consisted of examining the effects of ten ablations in the input data and network architecture that are commonly shared between CoRe++ and CoRe. The first ablation involved training CoRe++ with a single-channel depth image instead of the original four-channel RGB-D image. This ablation gave us a better understanding of the effect of adding color channels to the input data for completing the 3D shape. The second ablation involved training CoRe++ with an RGB-D image clipped to the bounding box instead of the original mask. This ablation gave us insight into how the final application on the potato harvester should look like: one based on an object detection model or one based on an instance segmentation model. Logically, this ablation also gave insight into the future annotation effort, which would be significantly higher when using an instance segmentation model. The third ablation involved reducing the encoder’s network architecture from seven to five convolutional blocks (layers 1-10 in Table 1 followed by a flatten layer and a fully connected layer). This ablation gave us insight into the effect of using an even lighter network in the situation of hardware constraints. The fourth ablation involved removing the pooling layers (the even layers in Table 1) to investigate their effect on the overall performance. The fifth ablation involved replacing CoRe++’s original LeakyReLU activation with the standard ReLU activation. The sixth ablation involved disabling the contrastive loss element in the loss function so that only the MSE loss was used during training. The seventh to the tenth ablation involved increasing or decreasing the learning rate in increments of two and five. This resulted in using a learning rate of 5$\cdot$10-4 (learning rate $\cdot$ 5), 2$\cdot$10-4 (learning rate $\cdot$ 2), 5$\cdot$10-5 (learning rate $\cdot$ 0.5) and 2$\cdot$10-5 (learning rate $\cdot$ 0.2).
@@ 3.1 The effect of the latent size on the 3D shape completion result
For both CoRe and CoRe++, the best latent size was 32, followed by 64. This result is consistent with our hypothesis that the medium-sized latent vectors are better for compressing the data into meaningful 3D representations. Figure 9 shows that latent sizes 32 and 64 have the most diverse 3D shapes and sizes after latent space interpolation. Compared to the other sizes, latent sizes 32 and 64 have both spherical and elongated 3D shapes, and the sizes show a proper sequential build-up. Latent size 8 and 256 are both worse in 3D shape completion performance (Table 3.1) and latent space interpolation (Figure 9). Latent size 8 could only produce five valid 3D shapes out of the seven interpolated latent vectors, while latent size 256 failed to produce a sequential size build-up. Latent sizes 16 and 256 produced mainly elongated 3D shapes, meaning that they had a limited ability to produce the more spherical shapes.
@@
Only for CoRe++ and latent sizes 32 and 64, the accuracy requirement was met as the corresponding RMSE values were less than 31.1 ml (this was the baseline value when using linear regression). The requirement for analysis speed was met for each latent size for both CoRe and CoRe++, as all of the analysis times were less than 16 ms (Table 3.1).
@@
Table 3: The effect of the latent size on the 3D shape completion results with DeepSDF, CoRe, and CoRe++. The upward arrows indicate the higher the better, and the downward arrows the lower the better. The values in bold are the best performing values per 3D shape completion method. The last two columns summarize whether the predefined accuracy and speed requirements were met. Note that the accuracy requirement was met if the RMSE on the volume was less than 31.1 ml (this was the RMSE when applying linear regression). The speed requirement was met if the analysis time was less than 16 ms. Refer to Section 2.4.1.
@@
Figure 9: 3D shape completion results for the seven latent space interpolations, visualized for each of the six tested latent sizes. The values above the 3D shapes are the volumes of the generated meshes in milliliters (ml).
Figure 10: (a) CoRe++’s best 3D shape completion result was achieved on a medium-sized potato tuber. (b) CoRe++’s worst 3D completion result was achieved on a large-sized and irregularly-shaped potato tuber. In both figure (a) and (b), the partial point cloud from the Realsense D405 is visualized on the left, and this point cloud needs to be completed. The second from the left visualizes the completion result with CoRe, including the volumetric estimate in milliliters. The colors of the 3D shape represent the Chamfer distances, where the orange and red colors are above 5 millimeters and the green and yellow colors below 5 millimeters. The second from the right visualizes the completion result with CoRe++. Most right is the 3D ground truth shape from the structure-from-motion method. Note that figure (a) and (b) were not at the same scale because they were zoomed differently to reveal better details.
Table 4 summarizes the effect of the potato size on the 3D shape completion result. There is a trend that the RMSE on the volumetric estimate is larger when the tuber is larger. A possible explanation is that the larger tubers are more concave (Table 4), meaning that they have more valleys and variations in their 3D curvature, making it harder to accurately reconstruct the 3D shape (especially when the concave parts are facing downward with respect to the camera’s perspective). The fact that larger Chamfer distances are observed on the largest tubers supports this explanation. For the relative volumetric errors, there is an opposite trend, meaning that the largest relative volumetric errors are observed on the smaller tubers.
@@
Table 5 summarizes the effect of the potato cultivar, and thus indirectly the shape of the potato tuber, on the 3D shape completion results. There is a trend that the Chamfer distance is lower when the tubers have a more spherical shape (i.e. elongation factors closer to 1.0). For the RMSE on the volumetric estimate, there is no clear trend regarding the tuber shape. What may explain this result is that the different cultivars had different tuber sizes: the cultivar Corolle had the relative smallest tubers and it was already shown in Table 4 that the RMSE is lowest when the tubers are smaller. The largest RMSE was observed on Sayaka, which was the cultivar with the largest tubers on average, but it was also the cultivar with the largest number of samples, meaning that there was probably more diversity in tuber shape and size.
@@
Figure 11 visualizes the RMSE on the volumetric estimate for the thirteen different image regions. The smallest RMSE of 18.2 ml was observed in the central horizontal region of the image between 350 and 400 pixels. The largest RMSE values were observed in the lower and upper parts of the image, indicating that these regions are not recommended for performing the 3D shape completion. These results may indicate that some form of lens distortion may have occurred in the peripheral regions of the camera’s field of view, leading to poorer 3D shape completion in these regions. There may also have been a higher degree of occlusion in these regions.
@@ 3.3 Ablation studies
The ablation study on CoRe++’s additions (Table 6) shows that the largest contribution to the overall performance was made by CoRe++’s validation method, which was based on GPU mesh generation and RMSE validation metric. Changing this validation method to the original validation method of CoRe with marching cubes mesh generation and Chamfer distance validation metric, led to the largest increase in both Chamfer distance (+32.1%) and RMSE on volumetric estimate (+122.1%). An obvious explanation for this result is that the final volumetric estimate benefits from having the RMSE optimized during training. Another explanation is that CoRe’s original marching cubes method seems unable to accurately reconstruct the 3D shapes of the potato tubers. This may be due to the chosen grid density of the marching cubes method which was optimized for high-throughput 3D reconstruction, but possibly came at the expense of the accuracy (something that can also be observed in Figure 10). Three other additions to CoRe++ that significantly improved the overall performance were the two data preprocessing steps (highlighted by the first and second ablation) and the loss function modification to MSE (highlighted by the fifth ablation).
Regarding the generic ablation study (Table 6), there are a few interesting outcomes. First, the choice of the activation function has a large impact on the overall performance. The fifth ablation shows that after replacing the LeakyReLU activation with a standard ReLU, the Chamfer distance increased by 46.4% and the RMSE by 96.0%. This outcome is consistent with that of Tomar (2022), who also found that ReLU under-performs compared to the more advanced variants of ReLU when testing an autoencoder on the Global Wheat Head dataset (David et al., 2020). In our case, we think that the LeakyReLU activation resulted in a better and faster network convergence during training, due to LeakyReLU’s ability to maintain non-zero gradients for negative inputs. A second interesting outcome was the relatively large impact of the RGB color channels on the overall performance. Our initial expectation was that only the depth image would contain relevant features for completing the 3D shape. There are two possible explanations for this result. First, the depth image has typically more noise than the RGB image, which may lead to less good feature extraction when only using the depth image. Second, the data augmentations for the RGB color channels were more extensive than for the depth channel, which may have resulted in a better generalization performance when using RGB-D images rather than just depth images. A third interesting outcome is that training without constrastive loss resulted in a lower Chamfer distance and a higher F-score. One explanation for this result is that the contrastive loss may only be beneficial if the images were obtained from different camera perspectives, as was the case in Magistri et al. (2022). In our experiment, the potato tubers were photographed from the same camera perspective, which resulted in relatively similar 3D shapes, making it more difficult for the contrastive loss function to separate the embedding space. Two other observations from the ablation study that are useful for implementing CoRe++ on an operational harvester are: an instance segmentation algorithm is preferred over an object detection algorithm (as highlighted by the second ablation), and a lighter encoder network may be more favorable for high-throughput shape completion, but it comes at the expense of the accuracy (as highlighted by the third ablation).
@@ 4 Discussion
Of the six latent sizes tested, we observed that a latent size of 32 outperformed the other sizes in terms of 3D shape completion and volumetric estimate. This suggests that a moderate latent size strikes a balance between representational capacity and model complexity, allowing RGB-D images to be encoded more effectively while not overfitting. The latter was clearly demonstrated by the latent space interpolation, which generated the most realistic potato shapes and sizes for latent size 32. Compared to the literature, our observations are similar to those of Ahmed and Longo (2022), who found that a latent size of 28 was best for optimizing a convolutional variational autoencoder on spectral topographic maps. Since Ahmed and Longo (2022) did their research on 25 latent sizes, this suggests that even better results could have been obtained if we had tested more latent sizes. Future work could investigate such a finer-grained latent size analysis or further explore the relationship between neural network architecture and latent size. Such an analysis could potentially provide new insights into how to further optimize the latent space for 3D shape completion tasks.
@@
With CoRe++ better 3D shape completion results were obtained compared to the linear regression model and the original CoRe implementation of Magistri et al. (2022). When we compare our results with those of Magistri et al. (2022), we can conclude that our results on potato are similar (in comparison with strawberry), or better (in comparison with sweet pepper). An important remark is that the obtained results may depend on the average complexity of the shape that has to be completed. Potato has on average a less complex shape than sweet pepper, which makes it easier for the network to learn a generic shape thus enabling a better performance. In future research, we want to test our CoRe++ model on crops or fruits with a more complex shape, such as pineapple, dragon fruit and Romanesco broccoli. We encourage fellow researchers to test our publicly available software on other 3D shapes within the agricultural domain or beyond.
@@
Our research has provided valuable insights and steps towards the practical application of CoRe++ on a potato harvester. Nevertheless, we think there are potential improvements in software and hardware that could further improve the performance. Regarding software, we would like to explore the use of other 3D shape completion networks, such as the one by Magistri et al. (2024), who used a transformer network. The obtained results with this new transformer network on sweet pepper and strawberry were significantly better than those of the original CoRe implementation (Magistri et al., 2022). Another advantage of using a transformer network is that it is end-to-end trainable, which is advantageous over our CoRe++ network that needs to be trained in two stages. A hardware improvement that could potentially improve the 3D shape completion is equipping the conveyor with rotating rollers in the area where the camera is placed. The rotating rollers cause the potato tubers to rotate gradually as they pass underneath the camera, meaning that almost the entire shape of the potato can be photographed. This can both simplify and improve the 3D shape completion. In a scenario like this, it would also be useful to test whether the contrastive loss function has an improving effect on completing the 3D shapes of the potato tubers when they rotate.
To further test the applicability of CoRe++ on a potato harvester, it is important to conduct a more in-depth evaluation on potato tubers of different cultivars. This evaluation will provide a better understanding of the overall generalization performance of CoRe++ at the farm level. In such an analysis, it is also important to conduct the test on potato tubers without the colored thumbtack, as the thumbtack’s color and shape can potentially affect the 3D shape completion result. A future evaluation could also benefit from performing the analysis on a more balanced dataset. In our study, potato tubers were randomly selected from the conveyor belt, but this led to an under-representation of the small and large tubers and those with irregular shape. In data-driven approaches such as ours, an unbalanced dataset would usually result in worse results on these minority cases, which is undesirable for a full-field harvest monitoring. Future research should also focus on reducing the degree of human subjectivity in selecting tubers on the conveyor belt by analyzing a batch of potatoes of different sizes at one time rather than manually selecting individual tubers. A final remark is that the performance of CoRe++ on a potato harvester will be affected by the instance segmentation algorithm that provides the necessary binary mask for CoRe++ to complete the 3D shape. Therefore, it is important to also investigate the sensitivity of CoRe++ to the segmentation outputs of the proposed instance segmentation algorithm.

### v3

@@ 2.3.3 Ablation studies
The second ablation study consisted of examining the effects of ten ablations in the input data and network architecture that were commonly shared between CoRe++ and CoRe. The first ablation involved training CoRe++ with a single-channel depth image instead of the original four-channel RGB-D image. This ablation gave us a better understanding of the effect of adding color channels to the input data for completing the 3D shape. The second ablation involved training CoRe++ with an RGB-D image clipped to the bounding box instead of the original mask. This ablation gave us insight into how the final application on the potato harvester should look like: one based on an object detection model or one based on an instance segmentation model. Logically, this ablation also gave insight into the future annotation effort, which would be significantly higher when using an instance segmentation model. The third ablation involved simplifying the encoder’s network architecture from seven to five convolutional blocks (layers 1-10 in Table 1 followed by a flatten layer and a fully connected layer). This ablation gave us insight into the effect on performance when using an even lighter network in the situation of hardware constraints. The fourth ablation involved removing the pooling layers (the even layers in Table 1) to investigate their effect on the overall performance. The fifth ablation involved replacing CoRe++’s original LeakyReLU activation with the standard ReLU activation. The sixth ablation involved disabling the contrastive loss element in the loss function so that only the MSE loss was used during training. The seventh to the tenth ablation involved increasing or decreasing the learning rate in increments of two and five. The resulting learning rates were 5$\cdot$10-4 (learning rate $\cdot$ 5), 2$\cdot$10-4 (learning rate $\cdot$ 2), 5$\cdot$10-5 (learning rate $\cdot$ 0.5) and 2$\cdot$10-5 (learning rate $\cdot$ 0.2).
2.4 Evaluation
For evaluating the 3D point cloud completion in our experiments, we used the Chamfer distance (Equation 1, Section 2.2.3) as performance metric. Besides the Chamfer distance, we also calculated the precision, recall and the F-score. The precision ($p$) was the percentage of reconstructed points within a certain distance ($d$) to the ground truth point cloud (Equation 6). As such, it represented the accuracy of the 3D shape completion. The recall ($r$) was the percentage of ground truth points within a certain distance ($d$) to the reconstructed point cloud (Equation 7). The recall represented the completeness of the 3D shape completion. For our evaluation, we used a distance ($d$) threshold of 5.0 mm. This threshold was adopted from Magistri et al. (2022), so that we could compare our results one-on-one with their results. After calculating the precision and recall, the F-score was obtained (Equation 8). The F-score was the harmonic mean between the precision and the recall, and it represented the percentage of 3D points that were correctly reconstructed.
[…]
For evaluating the volumetric estimate in our experiments, we used the RMSE metric, as presented in Equation 4. To evaluate the processing speed, we calculated the average 3D shape completion time on the test images. This processing speed analysis was performed on a Lenovo Legion Pro 7 16IRX8H laptop with an Intel i9-13900HX 2.2 GHz CPU with 32GB RAM and a NVIDIA GeForce RTX 4090 Laptop GPU with 16GB memory. This laptop was considered a suitable laptop for use on an operational harvester.
@@ 3.1 The effect of the latent size on the 3D shape completion result
For both CoRe and CoRe++, the best latent size was 32, followed by 64. This result is consistent with our hypothesis that the medium-sized latent vectors strike a better balance between compression and generalizability. Figure 9 shows that latent sizes 32 and 64 have the most diverse 3D shapes and sizes after latent space interpolation. Compared to the other sizes, latent sizes 32 and 64 have both spherical and elongated 3D shapes, and the sizes show a proper sequential build-up, indicating a better generalized latent space. Latent size 8 and 256 are both worse in 3D shape completion performance (Table 3.3) and latent space interpolation (Figure 9). Latent size 8 could only produce five valid 3D shapes out of the seven interpolated latent vectors, while latent size 256 failed to produce a sequential size build-up. Latent sizes 16 and 128 produced mainly elongated 3D shapes, meaning that they had a limited ability to produce the more spherical shapes.
@@
Only for CoRe++ and latent sizes 32 and 64, the accuracy requirement was met as the corresponding RMSE values were less than 31.1 ml (this was the baseline value when using linear regression). The requirement for analysis speed was met for each latent size for both CoRe and CoRe++, as all of the analysis times were less than 16 ms (Table 3.3). Regarding analysis time, no clear trend was observable when increasing or decreasing the latent vector size, implying that it does not matter for the analysis speed which size is chosen.
This experiment was conducted with CoRe++ and latent size 32, as this was the best performing combination in the previous experiment. Table 4 summarizes the effect of the potato size on the 3D shape completion result. There is a trend that the RMSE on the volumetric estimate is larger when the tuber is larger. A possible explanation is that the larger tubers are more concave (Table 4), meaning that they have more valleys and variations in their 3D curvature, making it harder to accurately reconstruct the 3D shape (especially when the concave parts are facing downward with respect to the camera’s perspective). It was demonstrated that especially the DeepSDF decoder produced large volumetric errors on tubers with a high concavity factor. This indicates that DeepSDF struggled in reconstructing the concave parts. The encoder produced the largest volumetric errors on the largest tubers in the test set (their volume was in between 225 and 357 ml), and this may be due to the under-representation of these tubers in our dataset (Figure 5). In-depth analysis on the five potato tubers with the largest 3D shape completion errors confirmed these trends, as these tubers had a high concavity factor (up to 0.45), were abnormally large or both. CoRe++’s largest error of 104 ml (which is visualized in Figure 9b), was on the second largest and the second most concave potato tuber in the test set. Thus, for a better generalization performance on the potato harvester, it may be beneficial to increase data diversity and expand the train set with more examples of large, concave tubers.
Table 5 summarizes the effect of the potato cultivar, and thus indirectly the shape of the potato tuber, on the 3D shape completion results. There is a trend that the Chamfer distance is lower when the tubers have a more spherical shape (i.e. elongation factors closer to 1.0). For the RMSE on the volumetric estimate, there is no clear trend regarding the tuber shape. What may explain this result is that the different cultivars had different tuber sizes: the cultivar Corolle had the relative smallest tubers and it was already shown in Table 4 that the RMSE is lowest when the tubers are smaller. The largest RMSE was observed on Sayaka, which was the cultivar with the largest tubers on average, but it was also the cultivar with the largest number of samples, meaning that there was probably more diversity in tuber shape and size. Thus, for better generalization performance on cultivars with highly varied tuber characteristics, it may be beneficial to extend the data augmentations when training the encoder and decoder.
Figure 11 visualizes the RMSE on the volumetric estimate for the thirteen different image regions. The smallest RMSE of 18.2 ml was observed in the central horizontal region of the image between 350 and 400 pixels. By performing the 3D shape completion only in this image region, the computational efficiency of the volumetric estimates can be improved in a practical setup. The largest RMSE values were observed in the lower and upper parts of the image, indicating that these regions are not recommended for performing the 3D shape completion. These results may also indicate that some form of lens distortion has occurred in the peripheral regions of the camera’s field of view, leading to poorer 3D shape completion in these regions. There may also have been a higher degree of occlusion in these regions.
3.3 Ablation studies
The ablation study on CoRe++’s additions (Table 6) shows that the largest contribution to the overall performance was made by CoRe++’s validation method, which was based on GPU mesh generation and RMSE validation metric. Changing this validation method back to the original validation method of CoRe with marching cubes mesh generation and Chamfer distance validation metric, led to the largest increase in both Chamfer distance (+32.1%) and RMSE on volumetric estimate (+122.1%). An obvious explanation for this result is that the final volumetric estimate benefits from having the RMSE optimized during training. Another explanation is that CoRe’s original marching cubes method seems unable to accurately reconstruct the 3D shapes of the potato tubers. This may be due to the chosen grid density of the marching cubes method which was optimized for high-throughput 3D reconstruction, but possibly came at the expense of the accuracy (something that can also be observed in Figure 10). Three other additions to CoRe++ that significantly improved the overall performance were the two data preprocessing steps (highlighted by the first and second ablation) and the loss function modification to MSE (highlighted by the fifth ablation).
Regarding the general ablation study (Table 6), there are a few interesting outcomes. First, the choice of the activation function had a large impact on the overall performance. The fifth ablation shows that after replacing the LeakyReLU activation with a standard ReLU, the Chamfer distance increased by 46.4% and the RMSE by 96.0%. This outcome is consistent with that of Tomar (2022), who also found that ReLU under-performs compared to the more advanced variants of ReLU when testing an autoencoder on the Global Wheat Head dataset (David et al., 2020). In our case, we think that the LeakyReLU activation resulted in a better and faster network convergence during training, due to LeakyReLU’s ability to maintain non-zero gradients for negative inputs. A second interesting outcome was the relatively large impact of the RGB color channels on the overall performance. Our initial expectation was that only the depth image would contain relevant features for completing the 3D shape. There are two possible explanations for this result. First, the depth image has typically more noise than the RGB image, which may lead to less good feature extraction when only using the depth image. Second, the data augmentations for the RGB color channels were more extensive than for the depth channel, which may have resulted in a better generalization performance when using RGB-D images rather than just depth images. A third interesting outcome is that training without constrastive loss resulted in a lower Chamfer distance and a higher F-score. One explanation for this result is that the contrastive loss may only be beneficial if the images were obtained from different camera perspectives, as was the case in Magistri et al. (2022). In our experiment, the potato tubers were photographed from the same top-view camera perspective, which resulted in relatively similar 3D shapes, making it more difficult for the contrastive loss function to separate the embedding space. Two other observations from the ablation study that are useful for implementing CoRe++ on an operational harvester are: an instance segmentation algorithm is preferred over an object detection algorithm (as highlighted by the second ablation), and a lighter encoder network may be more favorable for high-throughput shape completion, but it comes at the expense of the accuracy (as highlighted by the third ablation). In summary, the two ablation studies highlight that CoRe++’s validation method, data preprocessing, LeakyReLU activation, and inclusion of RGB channels were key to improving the 3D shape completion of potato tubers.
@@ 3.3 Ablation studies
Table 3: The effect of the latent size on the 3D shape completion results with DeepSDF, CoRe, and CoRe++. The upward arrows indicate the higher the better, and the downward arrows the lower the better. The values in bold are the best performing values per 3D shape completion method. Please refer to Section 2.3.1 for the accuracy and speed requirements and Section 2.4 for the performance metrics.
@@
Figure 9: 3D shape completion results for the seven latent space interpolations, visualized for each of the six tested latent sizes. The far left and far right meshes are the smallest and largest tubers in the test set between which the interpolation was performed.
Figure 10: (a) CoRe++’s best 3D shape completion result was achieved on a medium-sized potato tuber. (b) CoRe++’s worst 3D completion result was achieved on a large-sized and irregularly-shaped potato tuber. In (a) and (b), the partial point cloud is visualized on the left, the completion result from CoRe is visualized on the second from the left, and the one from CoRe++ is visualized on the second from the right. Most right is the 3D ground truth shape.
@@ 4 Discussion
Of the six latent sizes, we observed that a latent size of 32 outperformed the other sizes in terms of 3D shape completion and volumetric estimate. This suggests that a moderate latent size strikes a balance between representational capacity and model complexity, allowing RGB-D images to be encoded more effectively while not overfitting. The latter was clearly demonstrated by the latent space interpolation, which generated the most realistic potato shapes and sizes for latent size 32. Compared to the literature, our observations are similar to those of Ahmed and Longo (2022), who found that a latent size of 28 was best for optimizing a convolutional variational autoencoder on spectral topographic maps. Since Ahmed and Longo (2022) did their research on 25 latent sizes, this suggests that even better results could have been obtained if we had tested more latent sizes. In future work, a finer-grained latent size analysis could help identify optimal latent configurations for 3D shape completion tasks in agriculture, thereby enabling better estimates of crop shapes, volumes, and total yield.
@@
With CoRe++ better 3D shape completion results were obtained compared to the linear regression model and the original CoRe implementation of Magistri et al. (2022). When we compare our results with those of Magistri et al. (2022), we can conclude that our results on potato are similar (in comparison with strawberry), or better (in comparison with sweet pepper). An important remark is that the obtained results may depend on the average complexity of the shape that has to be completed. Shape complexity can impact model generalizability, affecting its ability to adapt to unseen crop instances. Potato has on average a less complex shape than sweet pepper, making it easier for the network to learn a generic shape enabling a better generalization performance. In future research, we want to test our CoRe++ model on crops or fruits with a more complex shape, such as pineapple, dragon fruit and Romanesco broccoli. We encourage fellow researchers to test our publicly available software on other 3D shapes within the agricultural domain or beyond.
@@
Our research has provided valuable insights and steps towards the practical application of CoRe++ on a potato harvester. Nevertheless, we think there are potential improvements in software and hardware that could further improve the performance. Regarding software, we would like to explore the use of other 3D shape completion networks, such as the one by Magistri et al. (2024), who used a transformer network. The obtained results with this new transformer network on sweet pepper and strawberry were significantly better than those of the original CoRe implementation (Magistri et al., 2022). An advantage of using a transformer network is that it is end-to-end trainable, enabling simultaneous optimization of all network weights, which is advantageous over our CoRe++ network that needs to be trained in two stages. A hardware improvement that could potentially improve the 3D shape completion is equipping the conveyor belt with rotating rollers in the area where the camera is placed. The rotating rollers cause the potato tubers to rotate gradually as they pass underneath the camera, meaning that almost the entire shape of the potato can be photographed. This can both simplify and improve the 3D shape completion. In a scenario like this, it would also be useful to test whether the contrastive loss function has an improving effect on completing the 3D shapes of the rotating potato tubers.
To further test the applicability of CoRe++ on a potato harvester, it is important to conduct a more in-depth evaluation on potato tubers of different cultivars. This evaluation will provide a better understanding of the overall generalization performance of CoRe++ at the farm level. In such an analysis, it is important to perform the test on potato tubers without the colored thumbtack, because the thumbtack’s color and shape may have influenced CoRe++’s extraction of distinguishable features during training. This could have unintentionally caused a stronger model trigger towards colors and shapes that do not appear in practical situations. A future evaluation could benefit from performing the analysis on a more balanced dataset. Our in-depth analysis revealed that the largest errors were on small and large tubers and those with irregular shape, and this might be problematic when analyzing an entire field. Future research should therefore focus on obtaining more balanced datasets or developing training tools that help the network better deal with under-represented samples.
When applying CoRe++ in practice, it is important to investigate its performance, scalability and interpretability as part of a larger system. In integrated systems, the performance of CoRe++ will be affected by the instance segmentation algorithm that provides the binary mask. It is therefore important to investigate the sensitivity of CoRe++ to the segmentation output of such instance segmentation algorithm. In larger integrated systems, there may be constraints towards the scalability of CoRe++. For instance, when using multi-row harvesters, there will be a larger throughput of potato tubers on the conveyor belt and this puts pressure on CoRe++’s ability to complete the 3D shape of all tubers on time. A straightforward solution is increasing the computational resources, but this will make the overall system more expensive. We believe it is better to develop a multi-threaded software approach, in which all 3D shape completion tasks are distributed across multiple concurrently-running threads. Regarding the system’s output, it is important to consider that the results must be post-processed in such a way that they can be easily interpreted by the end user. In its current setup, CoRe++ produces a volumetric estimation for each potato tuber, but these estimates have to be spatially summarized in such a way that a farmer can easily interpret the output. This means that in an integrated system, the outputs of CoRe++ have to be geo-referenced by means of a global navigation satellite system (GNSS). Such a GNSS-integrated system provides opportunities for integration with in-season crop growth data from drones, so that additional insights into the impact of management decisions can be obtained.